OpenStudy (anonymous):

Which of the following is NOT true of the fruit fly? A. One generation spans about 20 years. B. A new generation can be produced within weeks. C. There are 8 chromosomes in a diploid cell. D. Fruit flies are small and easy to maintain in a laboratory @Somy

OpenStudy (abhisar):

\(\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\\Huge\color{blue}{\mathfrak{~~~~Welcome~to~OpenStudy!~\ddot\smile}}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\)

OpenStudy (abhisar):

What do u think ?

OpenStudy (anonymous):

A? maybe

OpenStudy (abhisar):

BINGO !!

OpenStudy (abhisar):

Fruit flies or Drosophila Melanogester has a lifespan of mere 40-50 days.

OpenStudy (abhisar):

Got it ?

OpenStudy (anonymous):

Yep @Abhisar

OpenStudy (abhisar):

\(\Huge{\overset{\frown}{\normalsize \left( \begin{matrix} \Large\cdot \quad \cdot\\ \cdot\\ \huge \smile \end{matrix} \right)}}\normalsize \\ \;/\quad \;\;\quad \backslash\)

OpenStudy (somy):

deh @Abhisar this em human lol

OpenStudy (abhisar):

:D

OpenStudy (anonymous):

That is one of the reasons they are used as a test organism. Very cheap to keep around, no one really cares what you do to them, the have very short generation times and in a laboratory setting have consistent fecundity, and have a relatively small genome. The nematode C. elegans has the cell destiny of every cell from from the blastula (maybe earlier, cannot remember) to adult hood. Course it only has ~1050 cells as an adult.